A married couple has taxable income of $85,475. Determine their marginal tax rate and their average tax rate. (Round your answer to 2 decinal places)
Answer #1

Tax liability = 9086+(22%*(85475-78950)) = $10521.50

Marginal tax rate = 22%

Average tax rate = total tax/taxable income = 10521.50/85475 = 12.31%